Assume that the risk-free rate of interest is 4% and the expected rate of return on the market is 16%. A share of stock sells for $63 today. It will pay a dividend of $3 per share at the end of the year. Its beta is 1.1. What do investors expect the stock to sell for at the end of the year?

Expected Stock Price:
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

=4+1.1*(16-4)=17.2%

Required return=(D1/Current price)+Growth rate

0.172=(3/63)+Growth rate

Growth rate=0.172-(3/63)

=0.124380952

Expected price=63*(1+Growth rate)

=63*(1+0.124380952)

=$70.836